OpenStudy (unofficialllyy):

A copy machine makes 36 copies per minute. How many copies does it make in 4 minutes and 15 seconds?

OpenStudy (anonymous):

First convert the minutes into seconds, since they gave you minutes and seconds. 1 min = 60sec. 4:15 min = 255sec. So now, plot it. 36 copies in 60 sec. = x copies in 255 sec. Cross multiply: 36*255 = 9,180 & 60*x = 60x Divide: 9180/60 = x Solve: 151 = x